OpenStudy (anonymous):

ideal Gas Law question Lithium reacts with N2 according to the following reaction: 6Li (s) + N2 (g) -----> 2Li3N (s) Lithium is placed in an inflexible 1.00 L container with pure N2. The N2 pressure of 1.00 atm and a temperature of 298 K. One hour later the pressure of N2 has dropped to 0.920 atm and the temperature has dropped to 295K. Calculate the number of moles of N2 that reacted with the lithium. Help please

OpenStudy (aaronq):

You're finding the moles consumed through the temperature and pressure data. Use the ideal gas law, PV=nRT, Arrange it at the 2 instances, \(n_0=\dfrac{P_0V_0}{RT_0}\) \(n_1=\dfrac{P_1V_1}{RT_1}\) moles consumed: \(\Delta n=n_0-n_1=\dfrac{P_0V_0}{RT_0}-\dfrac{P_1V_1}{RT_1}\) R and V are constant, formula reduces to: \(\Delta n=n_0-n_1=\dfrac{V}{R}(\dfrac{P_0}{T_0}-\dfrac{P_1}{T_1})\)